### v2

# TACO-RL: Task Aware Prompt Compression Optimization with Reinforcement Learning

Abstract
The increasing prevalence of large language models (LLMs) such as GPT-4 in various applications has led to a surge in the size of prompts required for optimal performance, leading to challenges in computational efficiency. Prompt compression aims to reduce the inference cost by minimizing input tokens without compromising on the task performance. However, existing prompt compression techniques either rely on sub-optimal metrics such as information entropy or model it as a task-agnostic token classification problem that fails to capture task-specific information.
To address these issues, we propose a novel and efficient reinforcement learning (RL) based task-aware prompt compression method. To ensure low latency requirements, we leverage existing Transformer encoder-based token classification model while guiding the learning process with task-specific reward signals using lightweight REINFORCE algorithm. We evaluate the performance of our method on three diverse and challenging tasks including text summarization, question answering and code summarization. We demonstrate that our RL-guided compression method improves the task performance by 8% - 260% across these three scenarios over state-of-the-art compression techniques while satisfying the same compression rate and latency requirements.
TACO-RL: Task Aware Prompt Compression Optimization with Reinforcement Learning
1 Introduction
In recent years, Large Language Models (LLMs) have experienced a surge in popularity due to their impressive performance on a wide range of natural language processing tasks Brown et al. (2020); Chowdhery et al. (2023), ranging from question answering Ushio et al. (2023) to code generation Chen et al. (2021) to incident management Ahmed et al. (2023); Zhang et al. (2024); Goel et al. (2024). To effectively utilize these models, various prompting techniques have been introduced, such as In-Context Learning (ICL) Brown et al. (2020), Chain-of-Thought (CoT) Wei et al. (2023), and Retrieval Augmented Generation (RAG) Lewis et al. (2021). While these techniques improve the performance and efficacy of LLMs by providing them with relevant context and guidance, these lead to increase in input prompt context length which leads to higher inference cost and latency requirements.
To address this issue, several prompt compression techniques (that attempts to reduce the context length without loosing essential information) have been introduced. These existing work can be broadly categorized into two major threads: (a) Task-aware compression models Chuang et al. (2024), that generally finetune a task-specific decoder model that leads to high inference latency and cost; and (b) Task-agnostic compression models that either removes tokens/lexical units Jiang et al. (2023); Li et al. (2023) based on their information entropy or train a supervised binary token classification model using expert compressed examples Pan et al. (2024). Therefore, the existing solutions either fail to capture the task-specific behaviours or lead to high inference cost and latency. These challenges lead to two important research questions:
  • $\bullet$
    How can we design a prompt compression model that effectively leverages both bidirectional context and task-aware reward signals for better performance (Q1)?
  • $\bullet$
    To minimize inference latency and cost, how can we train a cheaper model with proper guidance from task-specific reward signals (Q2)?
Figure 1: Encoder fine-tuning with RL using task specific reward signals on a Q/A task. The RL-guided compression model is able to understand the specificity of the question and retains the relevant context in the compressed prompt.
To address Q2, we build our work on the foundation laid by LLMLingua-2 Pan et al. (2024) which trained a task-agnostic encoder-based transformer model in a supervised setting for binary classification of input tokens where the target compressed prompt is generated using expensive and efficient LLM model such as GPT-4 Achiam et al. (2023). We propose a novel approach, Task-Aware Prompt Compression Optimization with Reinforcement Learning (TACO-RL), to guide the finetuned generic encoder model with task-specific reward signal (to address Q1) using on-policy reinforcement learning technique. As shown in Figure 1, during the model alignment process, we generate the task output from both the original and compressed prompt, and compute the task-specific reward signal using the divergence between these two outputs. These reward signals are then used to update the base encoder model using on-policy REINFORCE algorithm Williams (1992).
To illustrate the efficacy of our proposed method against state-of-the-art (SoTA) prompt compression methods, we conducted extensive experiments on three diverse and challenging tasks on open-source benchmark datasets: (a) text summarization on MeetingBank dataset; (b) question-answering tasks on Squad dataset and (c) code summarization on CodeSearchNet dataset. The empirical results across these tasks demonstrate that our RL-guided prompt compression method can improve the task performance by 8% - 260% over LLMLingua-2 and other benchmark approaches while ensuring same compression rate and inference cost or latency. To that end, our key contributions are as follows:
  • $\bullet$
    We propose an efficient task-aware prompt compression model that leverages task specific reward signals to fine-tune base language models using on-policy RL (i.e., REINFORCE algorithm).
  • $\bullet$
    We introduce a latency-aware encoder based Transformer model for prompt compression that is aligned with task-specific objectives and leads to low inference cost.
  • $\bullet$
    We conduct extensive experiments on three diverse tasks to evaluate the effectiveness of our proposed method and demonstrate that our TACO-RL method provides significant improvements over SoTA methods. We will open-source the code upon publication.
2 Related Work
Prompt compression for LLMs.
Prompt compression shortens the input prompt to improve the inference efficiency of LLMs over long context. The form of the compressed prompts can vary, including token pruning, abstractive summarization, prompt paraphrasing, or soft prompt tuning. For example, token pruning trims less important tokens from the prompt (Li et al., 2023), while abstractive summarization or prompt paraphrasing aims to condense the semantics into shorter concise text Xu et al. (2024). Soft prompt tuning, on the other hand, converts the original prompt into a vector (Mu et al., 2024). Among these methods, token pruning has proven to be particularly effective due to its flexibility and smaller computational overhead compared to other methods. Prior work on prompt pruning, such as Selective Context (Li et al., 2023) and LLMLingua (Jiang et al., 2023, 2024), use heuristic metrics to compute token importance and trim less important tokens. LLMLingua-2 (Pan et al., 2024) trains a transformer-based classifier on compression data distilled from GPT-4 to decide whether to prune a token. While these task-agnostic prompt compression methods are effective and generalize to some tasks, they still struggle to model token importance in specific tasks or domains.
RL-based prompt compression.
RL-based methods have also been applied to prompt compression. For example, Jung and Kim (2024) leverage RL to train an MLP classifier conditioned on the task language model for token pruning. However, their compressor depends on the hidden representations of a white-box model, and the prompts they consider are usually short instructions rather than long contexts. Chuang et al. (2024) use RL to train a generative language model to compress prompts, which typically has a high computational overhead for compression. They consider classification tasks where there is a more straightforward signal for reward. Huang et al. (2024) apply context pruning to in-context learning (ICL) examples, training a hierarchical pruner with RL to select more relevant ICL examples and preserve more important tokens in the selected examples for better demonstrations in mathematical reasoning. In addition to the different design choices and compression goals, our work demonstrates how we effectively combine offline and online training to obtain a task-aware prompt compression model that can better model token importance with low inference cost and latency.
3 Background
In this section, we provide an overview of the prompt compression and reinforcement learning (RL) framework, which are used as building blocks for our method.
3.1 Prompt Compression
The goal of prompt compression is to reduce the context length without loosing the essential information of original prompt. To ensure low latency and inference cost, we build our work on recently introduced LLMLingua-2 Pan et al. (2024) framework that translate the problem into a binary token classification problem using an encoder-based Transformer model. Given an input prompt $X=(x_{1},x_{2},...,x_{N})$ with $N$ tokens, the compression process is defined as:
$\mathbf{H}=\text{Encoder}(X)\in\mathbb{R}^{d\times N}$
where $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},...,\mathbf{h}_{N}]$, and $\mathbf{h}_{i}\in\mathbb{R}^{d}$ is the encoded representation of token $x_{i}$. For each token, a binary classification probability is computed:
$\mathbf{p}_{i}=\text{softmax}(\mathbf{W}\mathbf{h}_{i}+\mathbf{b})$
where $\mathbf{W}\in\mathbb{R}^{2\times d}$ and $\mathbf{b}\in\mathbb{R}^{2}$ are learnable parameters. The compression decision $y_{i}$ for the i-th token is determined by:
$y_{i}=\begin{cases}1&\text{if }\mathbf{p}_{i}[1]\geq 0.5\\ 0&\text{otherwise}\end{cases}$
The compressed prompt $X_{c}$ is then constructed by retaining only the tokens where $y_{i}=1$. The compression effectiveness is computed by Compression Rate ($\tau=|X_{c}|/|X|$), while the Compression Ratio ($C.R.=1/\tau$) represents the factor by which the prompt is compressed.
3.2 REINFORCE Algorithm
Reinforcement learning (RL) has proven to be effective for domain alignment of language models. The seminal work of Ouyang et al. (2022) demonstrate that the learning of decoder based Transformer (e.g., GPT) models can be represented as bandit environment and off-policy RL algorithms such as Proximal Policy Optimization (PPO) Schulman et al. (2017) with task specific reward signals can be an effective tool for domain alignment. However, as off-policy RL is generally computationally expensive and sample inefficient, we primarily focus on guiding our compression model with on-policy RL method. The REINFORCE algorithm Williams (1992) is a popular on-policy policy gradient method that can be used to optimize a parameterized policy. Let $\pi_{\theta}(a|s)$ denote a parameterized policy with parameters $\theta$ that given a state $s$, can generate the probability of executing action $a$. Our goal is to optimize the parameters $\theta$ to maximize the expected return: $J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)]$, where $R(\tau)$ denotes the cumulative rewards obtained for trajectory $\tau$. The gradients used to update the policy parameters are computed using Eq. 1.
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[\nabla_{\theta}\log\pi_{\theta}(a|s)R(\tau)]$ (1)
4 Methodology
In this work, we combine the power of latency aware encoder-based Transformer model for token classification and on-policy RL algorithm for developing an efficient task-aware prompt compression model. We now present our proposed method Task-Aware Prompt Compression Optimization with Reinforcement Learning (TACO-RL).
4.1 TACO-RL
Our proposed TACO-RL framework has 3 key components: (1) A base encoder policy for action sampling and token classification, (2) Task specific reward signal computation; and (3) Policy optimization using on-policy RL method.
Encoder Model and Action Sampling.
Given an input prompt sequence $\mathbf{x}=(x_{1},\ldots,x_{n})$, our encoder policy predicts probabilities $\mathbf{p}=(p_{1},\ldots,p_{n})$, where $p_{i}=P(a_{i}=1\mid x_{i})$ represents the probability of preserving token $x_{i}$. We sample binary actions $\mathbf{a}=(a_{1},\ldots,a_{n})$ from these probabilities:
$a_{i}\sim\text{Bernoulli}(p_{i})\quad\forall i\in\{1,\ldots,n\}$ (2)
The compressed prompt $\mathbf{x}_{c}$ is then constructed by retaining only the tokens where $a_{i}=1$.
Reward Calculation.
We leverage an effective but relatively cheaper LLM model, GPT-3.5-Turbo model to generate outputs $y_{\text{orig}}$ and $y_{\text{comp}}$ from the original and compressed prompts, respectively. The reward $r$ is then computed based on task-specific metrics $\mathcal{M}$ (see Section 4.2):
$r=\begin{cases}\mathcal{M}(y_{\text{comp}},y_{\text{orig}}),&\text{if }-L\leq\delta<L\\ r_{0},&\text{otherwise}\end{cases}$ (3)
Here, $\mathcal{M}(y_{\text{comp}},y_{\text{orig}})$ denotes the divergence metric between the output of original and compressed prompts. $r_{0}$ is a negative constant reward for out-of-range compression. $\delta(=|\mathbf{x}_{c}|-c\cdot|\mathbf{x}|)$ denotes the divergence from expected compression rate. We use two important parameters to control the learning process, a compression flexibility controller $c$ and a tolerance threshold $L$.
Compression Flexibility Controller. The compression flexibility controller $c$ is a tunable hyperparameter that represents a baseline proportion for the number of tokens to be retained in the compressed prompt relative to the original prompt. A smaller value of $c$ enforces stricter compression (fewer tokens to be retained), while a larger value allows more tokens to be preserved. This provides flexibility in controlling the trade-off between output quality and inference cost.
Tolerance Threshold. As our policy is unconstrained, the action sampling may not always satisfy the compression ratio requirements during training process. Therefore, we define tolerance threshold parameter, $L$ to control the divergence from expected number of compressed tokens. Let $\delta$ denote the divergence between the actual and expected number of compressed tokens. To ensure a smooth learning process, we allow the divergence value $\delta$ to fall within the range $[-L,L]$ and if this criteria is met, then the task-specific positive reward signal is propagated. However, to ensure that the compressed prompt is neither excessively short nor unnecessarily long, a constant negative reward $r_{0}$ is applied to penalize extreme deviation. This mechanism stabilizes the compression process and guides the model towards generating prompts of desired lengths. It should be noted that during inference, we always satisfy the exact compression rate by sampling $c\cdot|x|$ tokens with highest probability.
Policy Optimization.
Finally, with the task-specific reward signals $r$, we update the encoder policy parameters using the REINFORCE Williams (1992) algorithm. The loss function is defined as:
$\mathcal{L}=-r\sum_{i=1}^{n}\log p(a_{i}\mid x_{i})-\lambda H(\mathbf{p})$ (4)
where $H(\mathbf{p})$ represents Shanon’s entropy regularization Shannon (1948) and $\lambda$ balances the tradeoff between reward and exploration. The gradient of the loss function with respect to the model parameters $\theta$ is calculated using Eq. 5:
$\begin{split}\nabla_{\theta}\mathcal{L}=-\mathbb{E}_{a\sim p(\cdot\mid x;\theta)}\left[r\nabla_{\theta}\log p(a_{i}\mid x_{i};\theta)\right]\\ -\lambda\nabla_{\theta}H(\mathbf{p})\end{split}$ (5)
By iteratively optimizing this objective, our approach refines the encoder’s ability to generate compressed prompts that retain essential task-specific information while minimizing prompt length.
Overall Approach.
Algorithm 1 delineates the key steps of TACO-RL. We begin with a task-agnostic encoder-based Transformer model Pan et al. (2024) for token classification. In each epoch, we generate the compressed prompt for every training data from the current encoder policy $\pi_{\theta}$ and compute the output of both original and compressed prompt using GPT-3.5-turbo model. Based on these outputs, we compute the task-specific output divergence metrics and use that as a positive reinforcement if the compression ratio requirements are met, otherwise we provide a small negative reward to penalize the constraint violation. Finally, using this reward signal we update the current policy parameter $\theta$ using Eq. 5. We execute this policy optimization process for $E$ epochs to generate the final optimized task-aware compression model $\pi^{*}_{\theta}$.
Input: Training set $\mathcal{D}$, initial encoder $\pi_{\theta}$, compression rate $c$, tolerance $L$, number of epochs $E$, Metric $\mathcal{M}$
Output: Optimized encoder policy $\pi_{\theta}^{*}$
for epoch $=1$ to $E$ do
       for $P\in\mathcal{D}$ do
             $\mathbf{H}\leftarrow\pi_{\theta}(P)$
             for $w_{i}\in P$ do
                   $p_{i}\leftarrow\text{softmax}(\mathbf{W}\mathbf{h}_{i}+\mathbf{b})$
                   $a_{i}\sim\text{Bernoulli}(p_{i})$
            $P_{c}\leftarrow\{w_{i}|a_{i}=1\}$
             $y_{\text{orig}},y_{\text{comp}}\leftarrow\text{GPT}(P),\text{GPT}(P_{c})$
             $\delta\leftarrow|P_{c}|-c|P|$
             $r\leftarrow\begin{cases}\text{$\displaystyle\mathcal{M}$}(y_{\text{comp}},y_{\text{orig}}),&\text{if }|\delta|\leq L\\ r_{0},&\text{otherwise}\end{cases}$
             $\mathcal{L}\leftarrow-r\sum_{i}\log p(a_{i}|w_{i})-\lambda H(\mathbf{p})$
             Compute $\nabla_{\theta}\mathcal{L}$ using Eq. 5
             $\theta\leftarrow\theta$ + $\nabla_{\theta}\mathcal{L}$ ;
return $\pi_{\theta}^{*}$
  // updated policy
Algorithm 1 TACO-RL
4.2 Task-specific Rewards
This section describes the task-specific reward formulations for summarization and question answering tasks.
4.2.1 Reward Formulation
For both task types, we generate outputs using the original prompt $\mathbf{x}$ and the compressed prompt $\mathbf{x}_{c}$ through a LLM (e.g., GPT-3.5-turbo):
$y_{\text{orig}}=\text{GPT}(\mathbf{x},[q]),y_{\text{comp}}=\text{GPT}(\mathbf{x}_{c},[q])$
where $q$ is the question for QA tasks.
The divergence metric $\mathcal{M}(y_{\text{comp}},y_{\text{orig}})$ in Equation 3 defines a generalized reward structure that can be applied to any task. However, the optimal reward signal may vary depending on the task. For example, for summarization, we would like to maximize the similarity between the two versions of summaries while minimizing hallucination in the summary from the compressed prompt. For question answering, the focus is on optimizing the accuracy and completeness of the generated answers.
Based on the generalized structure, we implement the following task-specific rewards:
4.2.2 Summarization Tasks
For summarization, we utilize the BLEU Papineni et al. (2002) score:
$\mathcal{M}_{\text{Sum}}=\text{BLEU}(y_{\text{comp}},y_{\text{orig}})$ (6)
As a precision-based verbatim similarity metric, BLEU effectively captures n-gram overlap between the summaries of the original and compressed contexts to promote content similarity, and minimizes introduction of inaccuracies and hallucinations.
4.2.3 Question Answering Task
For the QA task, we define the following precision and recall functions to measure the accuracy of the textual answer compared to the original, and we use the F1 score as the final divergence metric:
Precision =|$\displaystyle=\frac{|\{y_{i}|y_{i}\in y_{\text{orig}}\cap y_{\text{comp}}\}|}{|\{y_{i}|y_{i}\in y_{\text{comp}}\}|}$ (7)
Recall =|$\displaystyle=\frac{|\{y_{i}|y_{i}\in y_{\text{orig}}\cap y_{\text{comp}}\}|}{|\{y_{i}|y_{i}\in y_{\text{orig}}\}|}$ (8)
[…]
The F1 score balances precision and recall, helping to ensure that the compressed prompt retains essential context for accurate answer generation.
Models Bleu Rouge1 Rouge2 RougeL BertScore F1
0.50 (2x compression)
LLMLingua-2 - MeetingBank 18.68 54.20 29.45 40.14 90.69
LLMLingua-2 - Wikitext 16.71 (-1.97) 52.58 (-1.62) 27.73 (-1.72) 39.05 (-1.09) 90.47 (-0.21)
LLMLingua 5.90 (-12.78) 38.22 (-15.98) 14.02 (-15.42) 26.49 (-13.65) 87.65 (-3.04)
Selective Context 12.94 (-5.74) 46.30 (-7.90) 24.41 (-5.03) 34.56 (-5.58) 89.66 (-1.03)
TACO-RL (Ours) 21.35 (+2.67) 55.34 (+1.14) 31.88 (+2.43) 42.17 (+2.03) 90.95 (+0.26)
0.33 (3x compression)
LLMLingua-2 - MeetingBank 15.11 51.67 25.60 37.18 90.17
LLMLingua-2 - Wikitext 12.93 (-2.18) 49.38 (-2.29) 23.32 (-2.28) 35.34 (-1.83) 89.79 (-0.38)
LLMLingua 3.98 (-11.13) 32.62 (-19.05) 10.58 (-15.02) 22.58 (-14.60) 86.52 (-3.65)
Selective Context 8.80 (-6.31) 40.22 (-11.45) 19.28 (-6.32) 29.44 (-7.74) 88.67 (-1.50)
TACO-RL (Ours) 19.36 (+4.26) 53.67 (+1.99) 29.54 (+3.94) 40.01 (+2.83) 90.54 (+0.37)
0.25 (4x compression)
LLMLingua-2 - MeetingBank 12.80 49.40 22.77 34.77 89.78
LLMLingua-2 - Wikitext 10.98 (-1.82) 47.34 (-2.07) 20.68 (-2.09) 33.06 (-1.71) 89.34 (-0.44)
LLMLingua 3.51 (-9.29) 30.98 (-18.42) 9.64 (-13.13) 21.33 (-13.45) 86.20 (-3.58)
Selective Context 6.37 (-6.43) 36.22 (-13.18) 15.97 (-6.80) 26.07 (-8.71) 88.00 (-1.78)
TACO-RL (Ours) 17.61 (+4.81) 52.33 (+2.92) 27.84 (+5.07) 38.57 (+3.79) 90.26 (+0.48)
0.20 (5x compression)
LLMLingua-2 - MeetingBank 11.13 47.50 21.01 33.25 89.44
LLMLingua-2 - Wikitext 9.51 (-1.61) 45.33 (-2.17) 18.71 (-2.30) 31.27 (-1.98) 88.90 (-0.54)
LLMLingua 3.42 (-7.71) 30.53 (-16.98) 9.62 (-11.38) 21.15 (-12.09) 86.14 (-3.29)
Selective Context 4.82 (-6.31) 33.15 (-14.35) 13.55 (-7.46) 23.74 (-9.51) 87.49 (-1.94)
TACO-RL (Ours) 15.85 (+4.73) 50.56 (+3.06) 26.04 (+5.03) 36.81 (+3.56) 89.96 (+0.52)
0.166 (6x compression)
LLMLingua-2 - MeetingBank 9.80 45.82 19.19 31.64 89.12
LLMLingua-2 - Wikitext 8.55 (-1.25) 43.44 (-2.38) 17.06 (-2.13) 29.67 (-1.96) 88.53 (-0.59)
LLMLingua 3.19 (-6.62) 29.85 (-15.97) 9.47 (-9.72) 20.75 (-10.89) 86.07 (-3.05)
Selective Context 4.06 (-5.74) 31.34 (-14.48) 12.21 (-6.97) 22.31 (-9.33) 87.12 (-2.00)
TACO-RL (Ours) 14.25 (+4.45) 48.60 (+2.78) 24.51 (+5.33) 35.08 (+3.44) 89.68 (+0.56)
Results with Original Prompts 21.50 55.19 33.03 42.90 91.12
Table 1: Performance metrics for different models across various compression rates on the MeetingBank Dataset. Values in parentheses indicate deltas from the original LLMLingua-2 baseline.
Models Bleu Rouge1 RougeL
LLMLingua-2 - MeetingBank 18.68 54.20 40.14
TACO-RL - with Rouge1 19.89 54.40 41.22
TACO-RL - with RougeL 19.72 54.11 41.20
TACO-RL - with BLEU 21.35 55.34 42.17
Results with Original Prompts 21.50 55.19 42.90
Models QA F1 Score EM Score
LLMLingua-2 - Squad 62.70 38.02
TACO-RL - with Token Wise Score 56.67 35.14
TACO-RL - with F1 + Token Wise Score 68.03 44.64
TACO-RL - with F1 69.62 46.32
Results with Original Prompts 71.40 47.49
Table 2: Ablation study on the effect of different rewards on model performance at 2x compression rate on the MeetingBank dataset (left) and on the SQuAD dataset (right).
(2x)(3x)(4x)(5x)(6x)Compression Ratio00 01010 10102020 20203030 30304040 40405050 50506060 60607070 7070QA F1 Score1108974918866848204218131555141713232231184815451363125922461838154813911268273625922386219920671515 15152020 20202525 25253030 30303535 35354040 40404545 4545EM ScoreOriginal Context - F1 ScoreLLMLinguaSelective ContextLLMLingua-2 - WikitextLLMLingua-2 - SquadTACO-RL (ours)Original Context - EM Score
Figure 2: Comparison of QA F1 Scores and EM Counts across different compression rates for various models on the Squad Dataset. The bars represent QA F1 Scores, and the dashed lines represent EM Scores. The numbers on top of the bars represent the EM Counts. The two lines on top represent the scores with Original Context.
5 Experiments
To evaluate the effectiveness of our proposed approach, we conducted experiments on three tasks: text summarization, question answering, and code summarization.
5.1 Datasets
We use the following datasets:
MeetingBank Hu et al. (2023) - a text summarization dataset of meeting transcripts with $\sim$44k train examples, each of maximum 512 sequence length, and 862 test examples of various lengths.
SQuAD 2.0 Rajpurkar et al. (2016) - a question answering dataset where each example consists of a context and a corresponding question. Compression for the task is challenging because it requires preservation of different parts of the context depending on the question. We use a train subset of $\sim$34k samples and a test subset of 6k samples.
CodeSearchNet Husain et al. (2020) - a code dataset where one of the tasks is to generate summaries or docstring for Python code. Since the context is not in natural language, compression for this task presents a unique challenge. For this dataset we fine-tuned the encoder on a subset of $\sim$25k examples and then tested it on a subset of 1300 examples.
5.2 Base Model Training
Our training process involves two stages: (1) training a base model using a technique similar to LLMLingua-2 Pan et al. (2024), and (2) fine-tuning this base model using our novel reinforcement learning approach with task-specific rewards.
For each target dataset, we trained a base model on a distinct dataset with a similar distribution to ensure domain relevance while avoiding direct overlap. Specifically, for MeetingBank and SQuAD 2.0, we utilized the Wikitext dataset Merity et al. (2016) for base model training. In the case of CodeSearchNet Husain et al. (2020), we employed the Py150 dataset Raychev et al. (2016a). To train our base model, we followed the LLMLingua-2 approach and created annotated datasets using GPT-4. Across all experiments, the base models were consistently trained for 10 epochs and the values of other hyperparameters were set the same.
5.3 Experimental Setup
In our experiments, we employed the same architecture used in LLMLingua-2, xlm-roberta-large Conneau et al. (2020) with 561M parameters, as the backbone for our prompt compression technique. We replace the lm_head with a classification head on top. The experiments were performed on a compute instance equipped with 8 NVIDIA V100 GPUs (32 GB variants). For downstream evaluation of the compressed prompts, we utilized GPT-3.5-Turbo-1103 as our target language model. To ensure deterministic outputs and consistent results across multiple runs, we set the temperature to 0. During the fine-tuning phase, we employed a learning rate of 1e-6 in conjunction with a Cosine Annealing Loshchilov and Hutter (2017) scheduler to stabilize the training process. Further details regarding the experimental setup can be found in Appendix A.
(2x)(3x)(4x)(5x)(6x)Compression Ratio00 022 2244 4466 6688 881010 1010BLEU Score4.152.822.361.881.784.742.842.261.971.647.013.812.281.451.049.578.136.274.984.2211.239.878.187.026.01Original ContextLLMLingua-2 - Py150LLMLingua-2 - CodeSearchNetLLMLinguaTACO-RL (Rouge1)TACO-RL (Bleu)
Figure 3: Comparison of BLEU score across different compression rates for various models on the CodeSearchNet dataset.
5.4 Baselines
We use a LLMlingua-2 Pan et al. (2024) base model trained on the same datasets as the primary baseline for all our experiments. We also compare our method with two other state-of-the-art compression techniques: Selective Context Li et al. (2023) and LLMLingua Jiang et al. (2023).
5.5 Empirical Results
We conduct experiments on the three datasets over five different compression ratios. We observe that models trained at one choice of hyper parameters (see Table 3 in the appendix) generalize well over all compression ratios.
MeetingBank: Table 1 presents the results for the meeting text summarization task. We report BLEU, ROUGE, and BERT Score Zhang* et al. (2020). Our approach yields a significant performance improvement, with a performance boost of over 14% in BLEU scores at 2x compression, which further increases to 45% at 6x compression. We observe similar trends for other metrics. Notably, at 2x compression, our model performance is close to the results with using the original contexts.
SQuAD 2.0: Figure 2 shows the trend of the model performance with increasing compression ratios on the QA task. More detailed results are provided in Table 5 in the appendix. Our method outperforms the LLMLingua-2 baseline by 11% and 22% at 2x compression and 43% and 63% at 6x compression in F1 score and Exact Match score respectively.
CodeSearchNet: Figure 3 shows the results for the code summarization task. Despite the poor performance of the baseline models, our approach surpasses the results obtained with original prompts. Our method shows gains ranging from 1.3x at 2x compression to 2.6x at 6x compression in BLEU scores compared to the baseline. See Table 6 in the appendix for detailed results. A key reason for this improvement is that the compressor learns to retain only the parts of the context that docstrings tend to focus on, while the original prompts lead the downstream model to produce a more general summary.
5.6 Ablation Study
We conducted an ablation study to evaluate the impact of different divergence functions on model performance. For the summarization task, we compared the following metrics: ROUGE-1, ROUGE-L, and BLEU in Table 7. For QA, we tested three reward functions: F1 score, token-wise similarity, and a combination of both in Table 5. The token-wise similarity score is calculated based on token overlaps between the question and the context sentences. See Appendix C for a detailed description. Table 2 presents the effects of different rewards on model performance at the 2x compression rate.
6 Conclusion
In this paper, we introduce TACO-RL, a novel reinforcement learning-based prompt compression method to address the computational challenges in large language models (LLMs). Our approach leverages task-specific reward signals to fine-tune a Transformer encoder-based compression model using on-policy RL, enabling effective compression while maintaining performance. Our approach performs consistently well across high compression ratios (up to 6x) without requiring re-training. We demonstrate significant performance improvements of 63%, 45%, and up to 2.6x on the Text Summarization, QA, and Code Summarization tasks respectively, compared to strong baselines. Our work offers a promising approach for optimizing prompt compression, paving the way for more efficient and performant NLP systems.
Limitations
Our approach to prompt compression using reinforcement learning and encoder models has few limitations. Firstly, the fine-tuning process is sensitive to the choice of reward function, base model, and task-specific prompts. Experiments with different reward metrics, such as BLEU and Rouge in Tables 2, 6, 7, show that the choice of metric can significantly impact the performance of the encoder model. Additionally, the quality of the base model, which is trained on a similar dataset, can affect the fine-tuning results.
Secondly, the complexity of the task and size of the dataset play a crucial role in the fine-tuning process. While Figure 3 shows improved performance of the fine-tuned models on a code dataset, the performance increase drops significantly if the task is switched to code-completion from code summarization (See Appendix C). Large datasets are required for the model to learn the complexities effectively, especially when using reinforcement learning.
Finally, the fine-tuning process is computationally expensive, requiring significant resources and time. Optimizing the computational efficiency of the fine-tuning process is an important consideration for future work. Addressing these limitations will be crucial for improving the practicality and scalability of our approach.
Appendix A Details of Experiments
All the experiments were conducted on a 8 V100 cluster instance with 32 GB memory each.
Throughout all the experiments the input prompts were kept at 512 token length due to the fixed input sequence length of the encoder used. For training samples larger that this length, we chunked them into section of 512 tokens and used each chunk separately. For testing we compressed each chunk individually and then concatenated all compressed chunks into a final compressed prompts.
The time duration for the experiments is heavily dependent on the rate limits of the APIs employed, since at each step of training we need to call the API to get outputs from GPT3.5. Although for the Original Prompts the output can be obtained once and then reused for further epochs. We used API endpoints with max 300k TPM (Token Per Minute) limit. Despite the large TPM limit, the inference speed is greatly affected by traffic.
MeetingBank For the base model we used the WikiText dataset trained on roughly 23k examples for 10 epochs. This base model was then fine-tuned on $\sim$44k samples created from the MeetingBank Hu et al. (2023) after chunking, for 4 epochs using our approach. The test set comprised of 862 meeting transcripts.
For this dataset only, we used target summaries generated using GPT-4 instead of the using the original summaries. This was done to keep the experiments compatible with the LLMLingua Jiang et al. (2023); Pan et al. (2024) baselines.
SQuAD 2.0 The base model for this this dataset was the same as used in the MeetingBank dataset due to the similarity in distribution. In both datasets the context consists of general English language, making WikiText a good candidate for the base model.
For the fine-tuning, a subset of the original SQuAD 2.0 Rajpurkar et al. (2016) dataset with $\sim$34k samples was used with 15 epochs of training. This subset was created by removing very long and very short contexts. We also focused on having same contexts with different questions to test out approach more rigorously. The test set contained $\sim$6k examples.
CodeSearchNet We only used the Python subset of the entire corpus in our experiments. The base model for this dataset was trained on a subset of Py150 Raychev et al. (2016b) dataset with $\sim$20k samples for 10 epochs.
The fine-tuning comprised training on a subset of the CodeSearchNet Husain et al. (2020) dataset with $\sim$25k samples for 4 epochs. Since code contexts need to be summarized differently that text ones, this subset was created by removing very long and very short code contexts from the original corpus. This was done to allow the model to learn the important sections of the complete context rather than just tokens. We tested on a test subset with $\sim$1300 examples of python code context.
Table 3 shows the various hyper-parameter values used across the experiments:
Dataset Epochs LR c L $\lambda$ $r_{0}$
MeetingBank 4 $1e^{-6}$ 0.5 30 0.01 -0.1
SQuAD 2.0 15 $1e^{-6}$ 0.5 30 0.01 -0.1
CodeSearchNet 4 $1e^{-6}$ 0.5 30 0.001 -0.1
Table 3: Hyper-parameter choices for the experiments
0.330.50.66Compression Flexibility (c)0.100.10 0.100.100.120.12 0.120.120.140.14 0.140.140.160.16 0.160.160.180.18 0.180.180.200.20 0.200.20BLEU ScoreCompression Ratio2x3x4x5x6x
515304560Tolerance Threshold (L)0.120.12 0.120.120.140.14 0.140.140.160.16 0.160.160.180.18 0.180.180.200.20 0.200.20BLEU ScoreCompression Ratio2x3x4x5x6x
Figure 4: Impact on downstream task performance by different c, L values during training on the MeetingBank dataset.
Appendix B Impact of hyper-parameters on Training
The choice of hyper-parameters is heavily influenced by the datasets. Ours, being a RL based approach, requires tuning of these hyper-parameters for best results. We show the variation in results on the MeetingBank dataset as we tweak there parameters in Figure 4.
We notice that for certain values of the hyper-parameters the performance is maximum and then decreases on either side of the value. This effect is more dominant with values of the compression flexibility controller c. With L, there is a general trend of increase in performance with increasing hyper-parameter values.
Appendix C Code Completion Task
We also tried to use the idea of prompt compression for Python Code Completion task. Given a code context, we aimed to discard unnecessary tokens and used the compressed context to predict the last line of the context. Although we achieved improvement over the base model scores, the increase in performance was very subtle and also not of any use.
For the reward in this task, apart from the F1 score, we used a distance based reward that favoured the tokens close to the end of the context. This was done since the next line of the context depends greatly on the last few lines. We evaluated the performance on the same test set of CodeSearchNet dataset as previously used but with last line of the context removed. This is shown in Table 4.
Model QA F1 Best Sub. EM EM Count EM Score
Base Model - Py150 15.24 99.92 91.00 7.00
Ours 17.73 100.00 116.00 8.92
Table 4: Performance of our approach on Code Completion Task
Appendix D Token-wise Relevance Reward
For the SQuAD QA task, we explored a novel token-wise reward, distinct from the other single-value metric rewards. This approach assigns importance to individual tokens, with higher values indicating better preservation of task-relevant tokens in the compressed prompt. This is in contrast with the earlier reward formulation where a single metric value reward was used. The reward is formulated as follows:
Step 1: Sentence-level Similarity
Split the original context $\mathbf{x}$ into $m$ sentences, $\mathbf{x}=\{s_{1},s_{2},\ldots,s_{m}\}$. For each sentence $s_{i}$, compute its similarity with the question $q$:
$\text{sim}(s_{i},q)=\cos\left(\text{Emb}(s_{i}),\text{Emb}(q)\right)$ (10)
where $\text{Emb}(\cdot)$ denotes sentence embeddings from Sentence Transformers Reimers and Gurevych (2019).
Step 2: Token-level Similarity Assignment
Assign each token $x_{j}\in s_{i}$ its sentence’s similarity score:
$\text{sim}(x_{j})=\text{sim}(s_{i},q)\quad\text{for }x_{j}\in s_{i}$ (11)
Step 3: Binary Action Masking
Apply the binary action vector $\mathbf{a}=(a_{1},a_{2},\ldots,a_{n})$ to mask token-wise similarity scores:
$\mathbf{r}_{\text{masked}}=\mathbf{a}\odot\mathbf{r}_{\text{sim}}$ (12)
Step 4: Mean Token-wise Reward Calculation
Compute the mean of the masked vector for the final relevance reward $r_{\text{sim}}$:
$r_{\text{sim}}=\frac{1}{\sum a_{i}}\sum_{j=1}^{n}a_{j}\cdot\text{sim}(x_{j})$ (13)
Step 5: Reward Integration
The reward can be used standalone or combined with the F1 score:
$r_{\text{QA}}=r_{\text{F1}}+\alpha\cdot r_{\text{sim}}$ (14)
This combined reward aimed to balance accurate question answering with retention of contextually relevant tokens. However, contrary to expectations, this formulation did not improve performance and even degraded the base model’s results, as evidenced in Table 5. Moreover, when combined with the F1 score, it diminished the effectiveness of the F1 score as a standalone reward.
Models QA F1 Score Best Subspan EM EM Count EM Score
0.50 (2x compression)
LLMLingua-2 - Squad 62.70 99.83 2246 38.02
LLMLingua-2 - Wikitext 62.71 (+0.01) 99.86 (+0.03) 2231 (-15) 37.77 (-0.25)
LLMLingua 36.62 (-26.09) 98.87 (-0.96) 1108 (-1138) 18.76 (-19.27)
Selective Context 51.39 (-11.32) 98.68 (-1.15) 2042 (-204) 34.57 (-3.45)
TACO-RL - with Token Wise Score 56.67 (-6.03) 99.78 (-0.05) 2076 (-170) 35.14 (-2.88)
TACO-RL - with F1 + Token Wise Score 68.03 (+5.33) 99.90 (+0.07) 2637 (+391) 44.64 (+6.62)
TACO-RL - with F1 69.62 (+6.91) 99.92 (+0.08) 2736 (+490) 46.32 (+8.30)
0.33 (3x compression)
LLMLingua-2 - Squad 53.11 99.58 1838 31.12
LLMLingua-2 - Wikitext 52.86 (-0.25) 99.66 (+0.08) 1848 (+10) 31.28 (+0.17)
LLMLingua 30.57 (-22.54) 97.85 (-1.73) 974 (-864) 16.49 (-14.63)
Selective Context 46.37 (-6.75) 98.26 (-1.32) 1813 (-25) 30.69 (-0.42)
TACO-RL - with Token Wise Score 46.25 (-6.86) 99.31 (-0.27) 1636 (-202) 27.70 (-3.42)
TACO-RL - with F1 + Token Wise Score 62.79 (+9.68) 99.81 (+0.24) 2412 (+574) 40.83 (+9.72)
TACO-RL - with F1 65.84 (+12.73) 99.92 (+0.34) 2592 (+754) 43.88 (+12.76)
0.25 (4x compression)
LLMLingua-2 - Squad 46.37 99.27 1548 26.21
LLMLingua-2 - Wikitext 46.15 (-0.21) 99.61 (+0.34) 1545 (-3) 26.16 (-0.05)
LLMLingua 28.27 (-18.10) 97.58 (-1.69) 918 (-630) 15.54 (-10.67)
Selective Context 40.35 (-6.02) 97.90 (-1.37) 1555 (+7) 26.32 (+0.12)
TACO-RL - with Token Wise Score 39.85 (-6.51) 99.17 (-0.10) 1342 (-206) 22.72 (-3.49)
TACO-RL - with F1 + Token Wise Score 58.37 (+12.01) 99.71 (+0.44) 2195 (+647) 37.16 (+10.95)
TACO-RL - with F1 61.70 (+15.33) 99.88 (+0.61) 2386 (+838) 40.39 (+14.19)
0.20 (5x compression)
LLMLingua-2 - Squad 41.97 99.12 1391 23.55
LLMLingua-2 - Wikitext 41.49 (-0.48) 99.41 (+0.29) 1363 (-28) 23.07 (-0.47)
LLMLingua 26.96 (-15.01) 97.26 (-1.86) 866 (-525) 14.66 (-8.89)
Selective Context 37.37 (-4.61) 97.83 (-1.29) 1417 (+26) 23.99 (+0.44)
TACO-RL - with Token Wise Score 35.20 (-6.77) 98.97 (-0.15) 1167 (-224) 19.76 (-3.79)
TACO-RL - with F1 + Token Wise Score 54.57 (+12.59) 99.78 (+0.66) 2024 (+633) 34.26 (+10.72)
TACO-RL - with F1 57.92 (+15.95) 99.76 (+0.64) 2199 (+808) 37.23 (+13.68)
0.166 (6x compression)
LLMLingua-2 - Squad 38.74 98.98 1268 21.47
LLMLingua -2 - Wikitext 38.78 (+0.04) 99.27 (+0.29) 1259 (-9) 21.31 (-0.15)
LLMLingua 25.96 (-12.78) 97.14 (-1.84) 848.00 (-420) 14.36 (-7.11)
Selective Context 35.53 (-3.21) 97.82 (-1.17) 1323.00 (+55) 22.40 (+0.93)
TACO-RL - with Token Wise Score 31.98 (-6.76) 98.93 (-0.05) 1011 (-257) 17.12 (-4.35)
TACO-RL - with F1 + Token Wise Score 51.56 (+12.82) 99.64 (+0.66) 1867 (+599) 31.61 (10.14)
TACO-RL - with F1 55.46 (+16.72) 99.81 (+0.83) 2067 (+799) 34.99 (13.53)
Results with Original Prompts 71.40 99.93 2805 47.49
Table 5: Performance metrics for different models across various compression rates on the Squad Dataset. Values in parentheses indicate deltas from the LLMLingua-2 baseline model trained on Squad dataset.
Models Bleu Rouge1 Rouge2 RougeL BertScore F1
0.50 (2x compression)
LLMLingua-2 - CodeSearchNet 4.74 27.06 8.08 18.63 80.71
LLMLingua-2 - Py150 4.15 (-0.59) 27.10 (+0.04) 7.70 (-0.38) 18.38 (-0.25) 80.75 (+0.04)
LLMLingua 7.01 (+2.27) 31.35 (+4.29) 13.09 (+5.01) 23.35 (+4.72) 81.48 (+0.78)
Selective Context - - - - -
TACO-RL - with Rouge1 9.57 (+4.83) 34.97 (+7.91) 15.85 (+7.77) 26.20 (+7.57) 81.85 (+1.14)
TACO-RL - with Bleu 11.23 (+6.49) 35.66 (+8.60) 17.98 (+9.90) 27.36 (+8.73) 82.03 (+1.32)
0.33 (3x compression)
LLMLingua-2 - CodeSearchNet 2.84 23.72 5.54 16.08 80.25
LLMLingua-2 - Py150 2.82 (-0.02) 23.68 (-0.04) 5.24 (-0.30) 15.74 (-0.34) 80.27 (+0.02)
LLMLingua 3.81 (+0.97) 26.44 (+2.72) 8.27 (+2.72) 18.79 (+2.71) 80.72 (+0.46)
Selective Context - - - - -
TACO-RL - with Rouge1 8.13 (+5.29) 34.18 (+10.46) 14.55 (+9.01) 25.54 (+9.46) 81.80 (+1.55)
TACO-RL - with Bleu 9.87 (+7.03) 35.94 (+12.22) 17.97 (+12.43) 27.60 (+11.52) 82.18 (+1.93)
0.25 (4x compression)
LLMLingua-2 - CodeSearchNet 2.26 21.99 4.54 14.78 80.05
LLMLingua-2 - Py150 2.36 (+0.10) 22.37 (+0.38) 4.40 (-0.14) 14.86 (+0.08) 80.11 (+0.06)
LLMLingua 2.28 (+0.03) 23.64 (+1.65) 5.78 (+1.23) 16.41 (+1.64) 80.24 (+0.18)
Selective Context - - - - -
TACO-RL - with Rouge1 6.27 (+4.01) 33.19 (+11.20) 13.37 (+8.83) 24.58 (+9.80) 81.75 (+1.70)
TACO-RL - with Bleu 8.18 (+5.92) 35.49 (+13.50) 17.40 (+12.86) 27.43 (+12.65) 82.24 (+2.19)
0.20 (5x compression)
LLMLingua-2 - CodeSearchNet 1.97 21.01 3.99 14.21 79.96
LLMLingua-2 - Py150 1.88 (-0.09) 21.07 (+0.06) 3.83 (-0.16) 14.14 (-0.07) 79.98 (+0.02)
LLMLingua 1.45 (-0.52) 21.39 (+0.37) 4.27 (+0.29) 14.74 (+0.53) 79.94 (-0.02)
Selective Context - - - - -
TACO-RL - with Rouge1 4.98 (+3.01) 31.84 (+10.83) 12.03 (+8.04) 23.68 (+9.47) 81.63 (+1.67)
TACO-RL - with Bleu 7.02 (+5.05) 34.96 (+13.95) 16.90 (+12.91) 26.90 (+12.69) 82.16 (+2.20)
0.166 (6x compression)
LLMLingua-2 - CodeSearchNet 1.64 20.14 3.48 13.56 79.87
LLMLingua-2 - Py150 1.78 (+0.14) 20.69 (+0.55) 3.47 (-0.01) 13.95 (+0.39) 79.92 (+0.05)
LLMLingua 1.04 (-0.61) 19.93 (-0.21) 3.42 (-0.06) 13.70 (+0.14) 79.72 (-0.15)
Selective Context - - - - -
TACO-RL - with Rouge1 4.22 (+2.58) 30.96 (+10.82) 11.58 (+8.10) 23.00 (+9.44) 81.58 (+1.71)
TACO-RL - with Bleu 6.01 (+4.37) 34.31 (+14.17) 16.10 (+12.62) 26.36 (+12.80) 82.12 (+2.25)
Results with Original Prompts 9.80 33.83 15.79 25.39 81.69
Table 6: Performance metrics for different models across various compression rates on the CodeSearchNet Dataset. The scores for Selective Context are not added as it struggles to compress code data.
Models Bleu Rouge1 Rouge2 RougeL BertScore F1
0.50 (2x compression)
LLMLingua-2 - MeetingBank 18.68 54.20 29.45 40.14 90.69
Ours - with Rouge1 19.89 (+1.21) 54.40 (+0.20) 30.55 (+1.11) 41.22 (+1.08) 90.82 (+0.13)
Ours - with RougeL 19.72 (+1.04) 54.11 (-0.10) 30.56 (+1.12) 41.20 (+1.06) 90.79 (+0.10)
TACO-RL - with Bleu 21.35 (+2.67) 55.34 (+1.14) 31.88 (+2.43) 42.17 (+2.03) 90.95 (+0.26)
0.33 (3x compression)
LLMLingua-2 - MeetingBank 15.11 51.67 25.60 37.18 90.17
TACO-RL - with Rouge1 17.40 (+2.29) 52.50 (+0.83) 27.39 (+1.79) 38.75 (+1.57) 90.34 (+0.18)
TACO-RL - with RougeL 17.64 (+2.53) 52.48 (+0.81) 27.90 (+2.30) 39.03 (+1.85) 90.38 (+0.21)
TACO-RL - with Bleu 19.36 (+4.26) 53.67 (+1.99) 29.54 (+3.94) 40.01 (+2.83) 90.54 (+0.37)
0.25 (4x compression)
LLMLingua-2 - MeetingBank 12.80 49.40 22.77 34.77 89.78
TACO-RL - with Rouge1 15.73 (+2.93) 51.01 (+1.60) 25.53 (+2.76) 36.90 (+2.13) 90.02 (+0.24)
TACO-RL - with RougeL 15.68 (+2.88) 50.77 (+1.36) 25.82 (+3.05) 37.04 (+2.27) 90.03 (+0.25)
TACO-RL - with Bleu 17.61 (+4.81) 52.33 (+2.92) 27.84 (+5.07) 38.57 (+3.79) 90.26 (+0.48)
0.20 (5x compression)
LLMLingua-2 - MeetingBank 11.13 47.50 21.01 33.25 89.44
TACO-RL - with Rouge1 13.82 (+2.69) 49.19 (+1.68) 23.58 (+2.57) 35.27 (+2.03) 89.69 (+0.25)
TACO-RL - with RougeL 13.98 (+2.86) 48.73 (+1.23) 24.10 (+3.09) 35.35 (+2.10) 89.71 (+0.28)
TACO-RL - with Bleu 15.85 (+4.73) 50.56 (+3.06) 26.04 (+5.03) 36.81 (+3.56) 89.96 (+0.52)
0.166 (6x compression)
LLMLingua-2 - MeetingBank 9.80 45.82 19.19 31.64 89.12
TACO-RL - with Rouge1 12.78 (+2.98) 48.08 (+2.26) 22.49 (+3.30) 34.24 (+2.61) 89.49 (+0.36)
TACO-RL - with RougeL 13.22 (+3.42) 47.62 (+1.79) 22.83 (+3.64) 34.33 (+2.69) 89.50 (+0.38)
TACO-RL - with Bleu 14.25 (+4.45) 48.60 (+2.78) 24.51 (+5.33) 35.08 (+3.44) 89.68 (+0.56)
Results with Original Prompts 21.50 55.19 33.03 42.90 91.12
Table 7: Performance metrics for different models trained using different rewards across various compression rates on the MeetingBank Dataset. Values in parentheses indicate deltas from the original LLMLingua-2 baseline.

### v3

@@
 We evaluate the performance of our method on three diverse and challenging tasks including text summarization, question answering and code summarization.
-We demonstrate that our RL-guided compression method improves the task performance by 8% - 260% across these three scenarios over state-of-the-art compression techniques while satisfying the same compression rate and latency requirements.
+We demonstrate that our RL-guided compression method improves the task performance by 8% - 189% across these three scenarios over state-of-the-art compression techniques while satisfying the same compression rate and latency requirements.
 TACO-RL: Task Aware Prompt Compression Optimization with Reinforcement Learning
@@
 While these techniques improve the performance and efficacy of LLMs by providing them with relevant context and guidance, these lead to increase in input prompt context length which leads to higher inference cost and latency requirements.
-To address this issue, several prompt compression techniques (that attempts to reduce the context length without loosing essential information) have been introduced.
+To address this issue, several prompt compression techniques (which attempt to reduce the context length without losing essential information) have been introduced.
 These existing work can be broadly categorized into two major threads: (a) Task-aware compression models Chuang et al.
@@
 (2024).
-Therefore, the existing solutions either fail to capture the task-specific behaviours or lead to high inference cost and latency.
+Therefore, existing solutions either fail to capture task-specific behaviors or lead to high inference cost and latency.
 These challenges lead to two important research questions:
-  • $\bullet$
-    How can we design a prompt compression model that effectively leverages both bidirectional context and task-aware reward signals for better performance (Q1)?
-  • $\bullet$
-    To minimize inference latency and cost, how can we train a cheaper model with proper guidance from task-specific reward signals (Q2)?
+  • How can we design a prompt compression model that effectively leverages bidirectional context Devlin et al. (2019) and provides low inference latency (Q1)?
+  • To minimize the computational cost needed for adapting this model to a new task, how can we efficiently train a model with proper guidance from task-specific reward signals (Q2)?
 Figure 1: Encoder fine-tuning with RL using task specific reward signals on a Q/A task.
 The RL-guided compression model is able to understand the specificity of the question and retains the relevant context in the compressed prompt.
-To address Q2, we build our work on the foundation laid by LLMLingua-2 Pan et al.
-(2024) which trained a task-agnostic encoder-based transformer model in a supervised setting for binary classification of input tokens where the target compressed prompt is generated using expensive and efficient LLM model such as GPT-4 Achiam et al.
+To address Q1, we build our work on the foundation laid by LLMLingua-2 Pan et al.
+(2024) which trained a task-agnostic (while being inference latency-aware) encoder-based transformer model in a supervised setting for binary classification of input tokens where the target compressed prompt is generated using an expensive and efficient LLM model such as GPT-4 Achiam et al.
 (2023).
-We propose a novel approach, Task-Aware Prompt Compression Optimization with Reinforcement Learning (TACO-RL), to guide the finetuned generic encoder model with task-specific reward signal (to address Q1) using on-policy reinforcement learning technique.
+We propose a novel approach, Task-Aware Prompt Compression Optimization with Reinforcement Learning (TACO-RL), to guide the finetuned generic encoder model with task-specific reward signal (to address Q2) using on-policy reinforcement learning technique.
 As shown in Figure 1, during the model alignment process, we generate the task output from both the original and compressed prompt, and compute the task-specific reward signal using the divergence between these two outputs.
@@
 To illustrate the efficacy of our proposed method against state-of-the-art (SoTA) prompt compression methods, we conducted extensive experiments on three diverse and challenging tasks on open-source benchmark datasets: (a) text summarization on MeetingBank dataset; (b) question-answering tasks on Squad dataset and (c) code summarization on CodeSearchNet dataset.
-The empirical results across these tasks demonstrate that our RL-guided prompt compression method can improve the task performance by 8% - 260% over LLMLingua-2 and other benchmark approaches while ensuring same compression rate and inference cost or latency.
+The empirical results across these tasks demonstrate that our RL-guided prompt compression method can improve the task performance by 8% - 189% over LLMLingua-2 and other benchmark approaches while ensuring same compression rate and inference cost or latency.
 To that end, our key contributions are as follows:
-  • $\bullet$
+  • We introduce a latency-aware encoder based Transformer model for prompt compression that is aligned with task-specific objectives and leads to low inference cost.
   • We propose an efficient task-aware prompt compression model that leverages task specific reward signals to fine-tune base language models using on-policy RL (i.e., REINFORCE algorithm).
-  • $\bullet$
-    We introduce a latency-aware encoder based Transformer model for prompt compression that is aligned with task-specific objectives and leads to low inference cost.
-  • $\bullet$
   • We conduct extensive experiments on three diverse tasks to evaluate the effectiveness of our proposed method and demonstrate that our TACO-RL method provides significant improvements over SoTA methods.
@@
 3.1 Prompt Compression
-The goal of prompt compression is to reduce the context length without loosing the essential information of original prompt.
+The goal of prompt compression is to reduce the context length without losing the essential information of the original prompt.
 To ensure low latency and inference cost, we build our work on recently introduced LLMLingua-2 Pan et al.
-(2024) framework that translate the problem into a binary token classification problem using an encoder-based Transformer model.
+(2024) framework that translates the problem into a binary token classification problem using an encoder-based Transformer model.
+A key innovation in our approach is leveraging bidirectional context to make more informed compression decisions.
+Unlike unidirectional models, the encoder model captures contextual information from both left and right directions, enabling a more nuanced understanding of each token’s importance.
+This bidirectional context is crucial in determining the relevance of a token within the broader prompt context.
 Given an input prompt $X=(x_{1},x_{2},...,x_{N})$ with $N$ tokens, the compression process is defined as:
@@
 where $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},...,\mathbf{h}_{N}]$, and $\mathbf{h}_{i}\in\mathbb{R}^{d}$ is the encoded representation of token $x_{i}$.
+The contextual representation $\mathbf{h}_{i}$ captures the token’s significance by considering its interactions with surrounding tokens through self-attention mechanisms.
+This means the probability of preserving or removing a token is not determined in isolation, but by its relationship to the entire prompt context.
 For each token, a binary classification probability is computed:
@@
 $\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[\nabla_{\theta}\log\pi_{\theta}(a|s)R(\tau)]$ (1)
+Models Bleu Rouge1 Rouge2 RougeL BertScore F1
+0.50 (2x compression)
+LLMLingua-2 - MeetingBank 18.68 54.20 29.45 40.14 90.69
+LLMLingua-2 - Wikitext 16.71 (-1.97) 52.58 (-1.62) 27.73 (-1.72) 39.05 (-1.09) 90.47 (-0.21)
+LLMLingua 5.90 (-12.78) 38.22 (-15.98) 14.02 (-15.42) 26.49 (-13.65) 87.65 (-3.04)
+Selective Context 12.94 (-5.74) 46.30 (-7.90) 24.41 (-5.03) 34.56 (-5.58) 89.66 (-1.03)
+TACO-RL (Ours) 21.35 (+2.67) 55.34 (+1.14) 31.88 (+2.43) 42.17 (+2.03) 90.95 (+0.26)
+0.33 (3x compression)
+LLMLingua-2 - MeetingBank 15.11 51.67 25.60 37.18 90.17
+LLMLingua-2 - Wikitext 12.93 (-2.18) 49.38 (-2.29) 23.32 (-2.28) 35.34 (-1.83) 89.79 (-0.38)
+LLMLingua 3.98 (-11.13) 32.62 (-19.05) 10.58 (-15.02) 22.58 (-14.60) 86.52 (-3.65)
+Selective Context 8.80 (-6.31) 40.22 (-11.45) 19.28 (-6.32) 29.44 (-7.74) 88.67 (-1.50)
+TACO-RL (Ours) 19.36 (+4.26) 53.67 (+1.99) 29.54 (+3.94) 40.01 (+2.83) 90.54 (+0.37)
+0.25 (4x compression)
+LLMLingua-2 - MeetingBank 12.80 49.40 22.77 34.77 89.78
+LLMLingua-2 - Wikitext 10.98 (-1.82) 47.34 (-2.07) 20.68 (-2.09) 33.06 (-1.71) 89.34 (-0.44)
+LLMLingua 3.51 (-9.29) 30.98 (-18.42) 9.64 (-13.13) 21.33 (-13.45) 86.20 (-3.58)
+Selective Context 6.37 (-6.43) 36.22 (-13.18) 15.97 (-6.80) 26.07 (-8.71) 88.00 (-1.78)
+TACO-RL (Ours) 17.61 (+4.81) 52.33 (+2.92) 27.84 (+5.07) 38.57 (+3.79) 90.26 (+0.48)
+0.20 (5x compression)
+LLMLingua-2 - MeetingBank 11.13 47.50 21.01 33.25 89.44
+LLMLingua-2 - Wikitext 9.51 (-1.61) 45.33 (-2.17) 18.71 (-2.30) 31.27 (-1.98) 88.90 (-0.54)
+LLMLingua 3.42 (-7.71) 30.53 (-16.98) 9.62 (-11.38) 21.15 (-12.09) 86.14 (-3.29)
+Selective Context 4.82 (-6.31) 33.15 (-14.35) 13.55 (-7.46) 23.74 (-9.51) 87.49 (-1.94)
+TACO-RL (Ours) 15.85 (+4.73) 50.56 (+3.06) 26.04 (+5.03) 36.81 (+3.56) 89.96 (+0.52)
+0.166 (6x compression)
+LLMLingua-2 - MeetingBank 9.80 45.82 19.19 31.64 89.12
+LLMLingua-2 - Wikitext 8.55 (-1.25) 43.44 (-2.38) 17.06 (-2.13) 29.67 (-1.96) 88.53 (-0.59)
+LLMLingua 3.19 (-6.62) 29.85 (-15.97) 9.47 (-9.72) 20.75 (-10.89) 86.07 (-3.05)
+Selective Context 4.06 (-5.74) 31.34 (-14.48) 12.21 (-6.97) 22.31 (-9.33) 87.12 (-2.00)
+TACO-RL (Ours) 14.25 (+4.45) 48.60 (+2.78) 24.51 (+5.33) 35.08 (+3.44) 89.68 (+0.56)
+Results with Original Prompts 21.50 55.19 33.03 42.90 91.12
+Table 1: Performance metrics for different models across various compression rates ($\tau$) on the MeetingBank Dataset. Values in parentheses indicate deltas from the original LLMLingua-2 baseline.
 4 Methodology
@@
 $r_{0}$ is a negative constant reward for out-of-range compression.
-$\delta(=|\mathbf{x}_{c}|-c\cdot|\mathbf{x}|)$ denotes the divergence from expected compression rate.
+$\delta(=|\mathbf{x}_{c}|-c\cdot|\mathbf{x}|)$ denotes the divergence from expected compression.
 We use two important parameters to control the learning process, a compression flexibility controller $c$ and a tolerance threshold $L$.
@@
 This provides flexibility in controlling the trade-off between output quality and inference cost.
+Also, this parameter is only required to guide the encoder during training and is different from the compression ratio used during inference.
+We show that the a model trained with a single vale for $c$, generalizes to different compression ratios during training.
 Tolerance Threshold.
@@
 This mechanism stabilizes the compression process and guides the model towards generating prompts of desired lengths.
-It should be noted that during inference, we always satisfy the exact compression rate by sampling $c\cdot|x|$ tokens with highest probability.
+It should be noted that during inference, we always satisfy the exact compression rate by sampling $\tau\cdot|x|$ tokens with highest probability.
 Policy Optimization.
@@
 $\mathcal{L}=-r\sum_{i=1}^{n}\log p(a_{i}\mid x_{i})-\lambda H(\mathbf{p})$ (4)
-where $H(\mathbf{p})$ represents Shanon’s entropy regularization Shannon (1948) and $\lambda$ balances the tradeoff between reward and exploration.
+where $H(\mathbf{p})$ represents Shannon’s entropy regularization Shannon (1948) and $\lambda$ balances the tradeoff between reward and exploration.
 The gradient of the loss function with respect to the model parameters $\theta$ is calculated using Eq.
@@
 Overall Approach.
-Algorithm 1 delineates the key steps of TACO-RL.
+Algorithm 1 describes the key steps of TACO-RL.
 We begin with a task-agnostic encoder-based Transformer model Pan et al.
@@
 We execute this policy optimization process for $E$ epochs to generate the final optimized task-aware compression model $\pi^{*}_{\theta}$.
-Input: Training set $\mathcal{D}$, initial encoder $\pi_{\theta}$, compression rate $c$, tolerance $L$, number of epochs $E$, Metric $\mathcal{M}$
+Input: Training set $\mathcal{D}$, initial encoder $\pi_{\theta}$, compression controller $c$, tolerance $L$, number of epochs $E$, Metric $\mathcal{M}$
 Output: Optimized encoder policy $\pi_{\theta}^{*}$
@@
 As a precision-based verbatim similarity metric, BLEU effectively captures n-gram overlap between the summaries of the original and compressed contexts to promote content similarity, and minimizes introduction of inaccuracies and hallucinations.
+BLEU’s inclusion of a penalty for extra tokens helps prevent gaming of the reward system Skalse et al.
+(2022), unlike recall-based metrics such as Rouge which could be exploited by simply generating longer outputs.
+This design choice results in more stable training through better-calibrated rewards.
 4.2.3 Question Answering Task
@@
 The F1 score balances precision and recall, helping to ensure that the compressed prompt retains essential context for accurate answer generation.
-Models Bleu Rouge1 Rouge2 RougeL BertScore F1
-0.50 (2x compression)
-LLMLingua-2 - MeetingBank 18.68 54.20 29.45 40.14 90.69
-LLMLingua-2 - Wikitext 16.71 (-1.97) 52.58 (-1.62) 27.73 (-1.72) 39.05 (-1.09) 90.47 (-0.21)
-LLMLingua 5.90 (-12.78) 38.22 (-15.98) 14.02 (-15.42) 26.49 (-13.65) 87.65 (-3.04)
-Selective Context 12.94 (-5.74) 46.30 (-7.90) 24.41 (-5.03) 34.56 (-5.58) 89.66 (-1.03)
-TACO-RL (Ours) 21.35 (+2.67) 55.34 (+1.14) 31.88 (+2.43) 42.17 (+2.03) 90.95 (+0.26)
-0.33 (3x compression)
-LLMLingua-2 - MeetingBank 15.11 51.67 25.60 37.18 90.17
-LLMLingua-2 - Wikitext 12.93 (-2.18) 49.38 (-2.29) 23.32 (-2.28) 35.34 (-1.83) 89.79 (-0.38)
-LLMLingua 3.98 (-11.13) 32.62 (-19.05) 10.58 (-15.02) 22.58 (-14.60) 86.52 (-3.65)
-Selective Context 8.80 (-6.31) 40.22 (-11.45) 19.28 (-6.32) 29.44 (-7.74) 88.67 (-1.50)
-TACO-RL (Ours) 19.36 (+4.26) 53.67 (+1.99) 29.54 (+3.94) 40.01 (+2.83) 90.54 (+0.37)
-0.25 (4x compression)
-LLMLingua-2 - MeetingBank 12.80 49.40 22.77 34.77 89.78
-LLMLingua-2 - Wikitext 10.98 (-1.82) 47.34 (-2.07) 20.68 (-2.09) 33.06 (-1.71) 89.34 (-0.44)
-LLMLingua 3.51 (-9.29) 30.98 (-18.42) 9.64 (-13.13) 21.33 (-13.45) 86.20 (-3.58)
-Selective Context 6.37 (-6.43) 36.22 (-13.18) 15.97 (-6.80) 26.07 (-8.71) 88.00 (-1.78)
-TACO-RL (Ours) 17.61 (+4.81) 52.33 (+2.92) 27.84 (+5.07) 38.57 (+3.79) 90.26 (+0.48)
-0.20 (5x compression)
-LLMLingua-2 - MeetingBank 11.13 47.50 21.01 33.25 89.44
-LLMLingua-2 - Wikitext 9.51 (-1.61) 45.33 (-2.17) 18.71 (-2.30) 31.27 (-1.98) 88.90 (-0.54)
-LLMLingua 3.42 (-7.71) 30.53 (-16.98) 9.62 (-11.38) 21.15 (-12.09) 86.14 (-3.29)
-Selective Context 4.82 (-6.31) 33.15 (-14.35) 13.55 (-7.46) 23.74 (-9.51) 87.49 (-1.94)
-TACO-RL (Ours) 15.85 (+4.73) 50.56 (+3.06) 26.04 (+5.03) 36.81 (+3.56) 89.96 (+0.52)
-0.166 (6x compression)
-LLMLingua-2 - MeetingBank 9.80 45.82 19.19 31.64 89.12
-LLMLingua-2 - Wikitext 8.55 (-1.25) 43.44 (-2.38) 17.06 (-2.13) 29.67 (-1.96) 88.53 (-0.59)
-LLMLingua 3.19 (-6.62) 29.85 (-15.97) 9.47 (-9.72) 20.75 (-10.89) 86.07 (-3.05)
-Selective Context 4.06 (-5.74) 31.34 (-14.48) 12.21 (-6.97) 22.31 (-9.33) 87.12 (-2.00)
-TACO-RL (Ours) 14.25 (+4.45) 48.60 (+2.78) 24.51 (+5.33) 35.08 (+3.44) 89.68 (+0.56)
-Results with Original Prompts 21.50 55.19 33.03 42.90 91.12
-Table 1: Performance metrics for different models across various compression rates on the MeetingBank Dataset. Values in parentheses indicate deltas from the original LLMLingua-2 baseline.
 Models Bleu Rouge1 RougeL
@@
 Table 2: Ablation study on the effect of different rewards on model performance at 2x compression rate on the MeetingBank dataset (left) and on the SQuAD dataset (right).
-(2x)(3x)(4x)(5x)(6x)Compression Ratio00 01010 10102020 20203030 30304040 40405050 50506060 60607070 7070QA F1 Score1108974918866848204218131555141713232231184815451363125922461838154813911268273625922386219920671515 15152020 20202525 25253030 30303535 35354040 40404545 4545EM ScoreOriginal Context - F1 ScoreLLMLinguaSelective ContextLLMLingua-2 - WikitextLLMLingua-2 - SquadTACO-RL (ours)Original Context - EM Score
-Figure 2: Comparison of QA F1 Scores and EM Counts across different compression rates for various models on the Squad Dataset.
-The bars represent QA F1 Scores, and the dashed lines represent EM Scores.
-The numbers on top of the bars represent the EM Counts.
-The two lines on top represent the scores with Original Context.
 5 Experiments
-To evaluate the effectiveness of our proposed approach, we conducted experiments on three tasks: text summarization, question answering, and code summarization.
-5.1 Datasets
-We use the following datasets:
-MeetingBank Hu et al. (2023) - a text summarization dataset of meeting transcripts with $\sim$44k train examples, each of maximum 512 sequence length, and 862 test examples of various lengths.
+Our experimentation spans three different domains, each presenting unique compression challenges.
+5.1 Datasets and Task Diversity
+MeetingBank Hu et al.
+(2023) is a conversational summarization dataset with $\sim$44k train examples and 862 test samples.
+The dataset challenges our approach by requiring compression of complex, non-linear dialogue transcripts where contextual relevance depends on nuanced speaker interactions.
 SQuAD 2.0 Rajpurkar et al.
-(2016) - a question answering dataset where each example consists of a context and a corresponding question.
-Compression for the task is challenging because it requires preservation of different parts of the context depending on the question.
-We use a train subset of $\sim$34k samples and a test subset of 6k samples.
+(2016) presents a more sophisticated question-answering challenge.
+Unlike straightforward summarization, this dataset requires selective information extraction where compression must preserve only the most relevant context for a specific question.
+Using a representative subset of $\sim$34k training and 6k test examples, the dataset tests the model’s ability to perform context-aware compression.
+This task is particularly challenging because different questions may require preserving entirely different segments of the same context.
 CodeSearchNet Husain et al.
-(2020) - a code dataset where one of the tasks is to generate summaries or docstring for Python code.
-Since the context is not in natural language, compression for this task presents a unique challenge.
-For this dataset we fine-tuned the encoder on a subset of $\sim$25k examples and then tested it on a subset of 1300 examples.
+(2020) presents a unique non-natural language compression challenge in code documentation.
+From the original large-scale dataset, we curated a subset of $\sim$25k training and 1300 test examples focusing on Python code summarization.
+Unlike natural language, code compression demands preserving complex syntactic structures, algorithmic logic, and domain-specific semantic nuances.
+The challenge lies in distilling code context into meaningful summaries by identifying key functional components, understanding the overall structure and logic flow, and retaining critical variable names and comments that convey essential information for comprehensibility.
+By selecting datasets spanning summarization, question-answering, and technical documentation, we comprehensively assess our prompt compression technique.
+Each dataset represents a distinct real-world information compression scenario: conversational summarization, targeted information extraction, and technical context distillation.
 5.2 Base Model Training
@@
 5.3 Experimental Setup
+(2x)(3x)(4x)(5x)(6x)Compression Ratio00 01010 10102020 20203030 30304040 40405050 50506060 60607070 7070QA F1 Score1108974918866848204218131555141713232231184815451363125922461838154813911268273625922386219920671515 15152020 20202525 25253030 30303535 35354040 40404545 4545EM ScoreOriginal Context - F1 ScoreLLMLinguaSelective ContextLLMLingua-2 - WikitextLLMLingua-2 - SquadTACO-RL (ours)Original Context - EM Score
+Figure 2: Comparison of QA F1 Scores and EM Counts across different compression rates for various models on the Squad Dataset.
+The bars represent QA F1 Scores, and the dashed lines represent EM Scores.
+The numbers on top of the bars represent the EM Counts.
+The two lines on top represent the scores with Original Context.
 In our experiments, we employed the same architecture used in LLMLingua-2, xlm-roberta-large Conneau et al.
@@
 For downstream evaluation of the compressed prompts, we utilized GPT-3.5-Turbo-1103 as our target language model.
-To ensure deterministic outputs and consistent results across multiple runs, we set the temperature to 0.
+We fixed the temperature parameter at zero to guarantee reproducible outcomes and uniform results in our experiments.
 During the fine-tuning phase, we employed a learning rate of 1e-6 in conjunction with a Cosine Annealing Loshchilov and Hutter (2017) scheduler to stabilize the training process.
 Further details regarding the experimental setup can be found in Appendix A.
-(2x)(3x)(4x)(5x)(6x)Compression Ratio00 022 2244 4466 6688 881010 1010BLEU Score4.152.822.361.881.784.742.842.261.971.647.013.812.281.451.049.578.136.274.984.2211.239.878.187.026.01Original ContextLLMLingua-2 - Py150LLMLingua-2 - CodeSearchNetLLMLinguaTACO-RL (Rouge1)TACO-RL (Bleu)
-Figure 3: Comparison of BLEU score across different compression rates for various models on the CodeSearchNet dataset.
 5.4 Baselines
@@
 We observe that models trained at one choice of hyper parameters (see Table 3 in the appendix) generalize well over all compression ratios.
+We also conduct statistical significance tests on MeetingBank dataset results (see Table 9) to support our performance enhancement claims.
+(2x)(3x)(4x)(5x)(6x)Compression Ratio00 055 551010 10101515 15152020 20202525 25253030 30303535 3535BLEU Score18.3011.988.937.316.5419.9012.119.357.766.7923.5813.349.066.575.3129.9722.7717.5914.1512.3835.0428.8124.2321.1318.46LLMLingua-2 - CodeSearchNetLLMLingua-2 - Py150LLMLinguaTACO-RL (Rouge1)TACO-RL (Bleu)
+Figure 3: Comparison of BLEU score across different compression rates for various models on the CodeSearchNet dataset.
 MeetingBank: Table 1 presents the results for the meeting text summarization task.
@@
 CodeSearchNet: Figure 3 shows the results for the code summarization task.
-Despite the poor performance of the baseline models, our approach surpasses the results obtained with original prompts.
-Our method shows gains ranging from 1.3x at 2x compression to 2.6x at 6x compression in BLEU scores compared to the baseline.
+Even though the base model performs poorly, leveraging our approach, the final fine-tuned model significantly outperforms all other baselines.
+Our method shows gains ranging from 0.91x (91%) at 2x compression to 1.89x at 5x compression in BLEU scores compared to the LLMLingua-2 baseline.
 See Table 6 in the appendix for detailed results.
-A key reason for this improvement is that the compressor learns to retain only the parts of the context that docstrings tend to focus on, while the original prompts lead the downstream model to produce a more general summary.
+Additionally, Appendix E highlights the differences in the compression mechanisms between TACO-RL and LLMLingua-2.
 5.6 Ablation Study
-We conducted an ablation study to evaluate the impact of different divergence functions on model performance.
-For the summarization task, we compared the following metrics: ROUGE-1, ROUGE-L, and BLEU in Table 7.
-For QA, we tested three reward functions: F1 score, token-wise similarity, and a combination of both in Table 5.
-The token-wise similarity score is calculated based on token overlaps between the question and the context sentences.
-See Appendix C for a detailed description.
-Table 2 presents the effects of different rewards on model performance at the 2x compression rate.
+We performed an ablation study to assess the impact of different reward metrics on model performance.
+For the text summarization task, we compared the following metrics: ROUGE-1, ROUGE-L, and BLEU, as shown in Table 2.
+For a detailed comparison over various compression ratios, see Table 7.
+For the QA task, we evaluated three reward functions: F1 score, token-wise similarity, and a combination of both, as shown in Table 2 (with additional details in Table 5).
+The token-wise similarity score is based on token overlaps between the question and context sentences, with a detailed explanation provided in Appendix C.
+Additionally, we investigate the effect of the entropy regularization term $H(\mathbf{p})$ on the fine-tuned model’s performance on the MeetingBank dataset, as detailed in Table 9.
 6 Conclusion
-In this paper, we introduce TACO-RL, a novel reinforcement learning-based prompt compression method to address the computational challenges in large language models (LLMs).
+In this paper, we introduce TACO-RL, a novel reinforcement learning-based prompt compression method to address computational challenges in large language models (LLMs).
 Our approach leverages task-specific reward signals to fine-tune a Transformer encoder-based compression model using on-policy RL, enabling effective compression while maintaining performance.
 Our approach performs consistently well across high compression ratios (up to 6x) without requiring re-training.
-We demonstrate significant performance improvements of 63%, 45%, and up to 2.6x on the Text Summarization, QA, and Code Summarization tasks respectively, compared to strong baselines.
-Our work offers a promising approach for optimizing prompt compression, paving the way for more efficient and performant NLP systems.
+We demonstrate significant performance improvements of 45%, 63%, and up to 1.89x on the Text Summarization, QA, and Code Summarization tasks respectively, compared to strong baselines.
+Our work offers a promising approach to optimizing prompt compression, paving the way for more efficient and performant NLP systems.
 Limitations
@@
 Experiments with different reward metrics, such as BLEU and Rouge in Tables 2, 6, 7, show that the choice of metric can significantly impact the performance of the encoder model.
-Additionally, the quality of the base model, which is trained on a similar dataset, can affect the fine-tuning results.
-Secondly, the complexity of the task and size of the dataset play a crucial role in the fine-tuning process.
-While Figure 3 shows improved performance of the fine-tuned models on a code dataset, the performance increase drops significantly if the task is switched to code-completion from code summarization (See Appendix C).
-Large datasets are required for the model to learn the complexities effectively, especially when using reinforcement learning.
+Additionally, the quality of the base model, which is trained on a similar dataset, can significantly affect the fine-tuning results.
+Secondly, the intricacies of task complexity and dataset magnitude significantly influence the fine-tuning methodology.
+As illustrated in Figure 3, fine-tuned models demonstrate enhanced performance on code datasets, yet this improvement diminishes markedly when transitioning from code summarization to code-completion tasks (See Appendix C).
+Extensive documentation strings within code contexts can potentially skew the compression dynamics, as models might develop strategies to retain doc-string tokens as a form of reward manipulation.
+Moreover, comprehensive and expansive datasets are essential for effectively capturing and learning the nuanced complexities, particularly in reinforcement learning scenarios.
 Finally, the fine-tuning process is computationally expensive, requiring significant resources and time.
@@
 The test set comprised of 862 meeting transcripts.
-For this dataset only, we used target summaries generated using GPT-4 instead of the using the original summaries.
+For this dataset, we used target (reference) summaries generated using GPT-4 instead of the using the original summaries.
 This was done to keep the experiments compatible with the LLMLingua Jiang et al.
@@
 The test set contained $\sim$6k examples.
+During evaluation, original answers present in the dataset were treated as reference answers.
 CodeSearchNet We only used the Python subset of the entire corpus in our experiments.
@@
 (2016b) dataset with $\sim$20k samples for 10 epochs.
-The fine-tuning comprised training on a subset of the CodeSearchNet Husain et al.
-(2020) dataset with $\sim$25k samples for 4 epochs.
-Since code contexts need to be summarized differently that text ones, this subset was created by removing very long and very short code contexts from the original corpus.
-This was done to allow the model to learn the important sections of the complete context rather than just tokens.
-We tested on a test subset with $\sim$1300 examples of python code context.
+We fine-tuned our model on a curated subset of $\sim$25k samples from the CodeSearchNet Husain et al.
+(2020) dataset over 4 epochs.
+Recognizing the unique challenges in code context summarization, we preprocessed the dataset by eliminating extremely long and short code contexts.
+This strategic filtering enables the model to focus on learning the most informative sections of code contexts, rather than getting distracted by peripheral tokens or overly verbose implementations.
+The evaluation was conducted on a test subset comprising $\sim$1300 Python code examples.
+Due to the lack of usable summaries in the dataset, reference summaries were generated using complete original prompts with GPT 3.5.
 Table 3 shows the various hyper-parameter values used across the experiments:
@@
 Moreover, when combined with the F1 score, it diminished the effectiveness of the F1 score as a standalone reward.
+Appendix E Prompt Compression Comparison For Code Summarization
+E.1 Original Code Context
+1def yixia_download(url, output_dir = ’.’, merge = True, info_only = False, **kwargs):
+2 """wrapper"""
+3 hostname = urlparse(url).hostname
+4 if ’n.miaopai.com’ == hostname:
+5 smid = match1(url, r’n\.miaopai\.com/media/([^.]+)’)
+6 miaopai_download_by_smid(smid, output_dir, merge, info_only)
+7 return
+8 elif ’miaopai.com’ in hostname: #Miaopai
+9 yixia_download_by_scid = yixia_miaopai_download
+10 site_info = "Yixia Miaopai"
+11 scid = match1(url, r’miaopai\.com/show/channel/([^.]+)\.htm’) or \
+12 match1(url, r’miaopai\.com/show/([^.]+)\.htm’) or \
+13 match1(url, r’m\.miaopai\.com/show/channel/([^.]+)\.htm’) or \
+14 match1(url, r’m\.miaopai\.com/show/channel/([^.]+)’)
+15 elif ’xiaokaxiu.com’ in hostname: #Xiaokaxiu
+16 yixia_download_by_scid = yixia_xiaokaxiu_download
+17 site_info = "Yixia Xiaokaxiu"
+18 if re.match(r’http://v.xiaokaxiu.com/v/.+\.html’, url): #PC
+19 scid = match1(url, r’http://v.xiaokaxiu.com/v/(.+)\.html’)
+20 elif re.match(r’http://m.xiaokaxiu.com/m/.+\.html’, url): #Mobile
+21 scid = match1(url, r’http://m.xiaokaxiu.com/m/(.+)\.html’)
+22 else:
+23 pass
+Listing 1: Original Code Context
+E.2 Compressed Context - LLMLingua-2
+1 ’n.miaopai
+2(url[^’
+3 ’miaopai.com’
+4 Miaopai
+5 match1(url.com/show/channel/([^.]+)\.htm
+6 match1[^
+7 match1([^.]+)’)
+8 ’xiaokaxiu.com’
+9_download
+10 "Yixia Xiaokaxiu"
+11 re.match(r’http://v.xiaokaxiu.com/v/.+\.html’,
+12 match1(url r’http://v.xiaokaxiu/v/(.+).html’)
+13 re.match(r’http://m.xiaokaxiu.com/m/.+\.html’
+14 match1(url, r’http://m.xiaokaxiu.com/m(
+15 else pass’
+Listing 2: Compressed Code by LLMLingua-2
+E.3 Compressed Context - TACO-RL
+1def yixia_download(url, output_dir = ’.’, merge = True, info_only = False, **kwargs):
+2 """wrapper"""
+3 hostname =
+4 if ’n.miaopai.com’ == hostname:
+5 smid =
+6 return
+7 elif ’miaopai.com’ in hostname: #Miaopai
+8 "Yixia Miaopai"
+9 scid
+10 elif ’xiaokaxiu.com’
+11 Xiaokaxiu
+12 if re #PC
+13 scid
+14 elif
+15+
+16 pass
+Listing 3: Compressed Code by TACO-RL
 Models QA F1 Score Best Subspan EM EM Count EM Score
@@
 0.50 (2x compression)
-LLMLingua-2 - CodeSearchNet 4.74 27.06 8.08 18.63 80.71
-LLMLingua-2 - Py150 4.15 (-0.59) 27.10 (+0.04) 7.70 (-0.38) 18.38 (-0.25) 80.75 (+0.04)
-LLMLingua 7.01 (+2.27) 31.35 (+4.29) 13.09 (+5.01) 23.35 (+4.72) 81.48 (+0.78)
+LLMLingua-2 - CodeSearchNet 18.30 51.57 22.50 37.94 90.27
+LLMLingua-2 - Py150 19.90 (+1.60) 52.93 (+1.36) 24.35 (+1.85) 39.65 (+1.71) 90.61 (+0.34)
+LLMLingua 23.58(+5.28) 55.23 (+3.66) 28.64 (+6.14) 43.41 (+5.47) 91.12 (+0.85)
 Selective Context - - - - -
-TACO-RL - with Rouge1 9.57 (+4.83) 34.97 (+7.91) 15.85 (+7.77) 26.20 (+7.57) 81.85 (+1.14)
-TACO-RL - with Bleu 11.23 (+6.49) 35.66 (+8.60) 17.98 (+9.90) 27.36 (+8.73) 82.03 (+1.32)
+TACO-RL - with Rouge1 29.97 (+11.67) 59.32 (+7.75) 33.99 (+11.49) 47.45 (+9.51) 91.76 (+1.49)
+TACO-RL - with Bleu 35.04 (+16.74) 61.26 (+9.69) 38.83 (+16.34) 50.77 (+12.84) 92.26 (+1.98)
 0.33 (3x compression)
-LLMLingua-2 - CodeSearchNet 2.84 23.72 5.54 16.08 80.25
-LLMLingua-2 - Py150 2.82 (-0.02) 23.68 (-0.04) 5.24 (-0.30) 15.74 (-0.34) 80.27 (+0.02)
-LLMLingua 3.81 (+0.97) 26.44 (+2.72) 8.27 (+2.72) 18.79 (+2.71) 80.72 (+0.46)
+LLMLingua-2 - CodeSearchNet 11.98 44.95 15.83 32.07 89.06
+LLMLingua-2 - Py150 12.11 (+0.14) 44.94 (-0.01) 15.81 (-0.02) 32.10 (+0.02) 89.10 (+0.04)
+LLMLingua 13.40 (+1.42) 46.29 (+1.34) 18.21 (+2.40) 34.49 (+2.42) 89.33 (+0.28)
 Selective Context - - - - -
-TACO-RL - with Rouge1 8.13 (+5.29) 34.18 (+10.46) 14.55 (+9.01) 25.54 (+9.46) 81.80 (+1.55)
-TACO-RL - with Bleu 9.87 (+7.03) 35.94 (+12.22) 17.97 (+12.43) 27.60 (+11.52) 82.18 (+1.93)
+TACO-RL - with Rouge1 22.77 (+10.79) 54.39 (+9.44) 27.34 (+11.50) 41.81 (+9.74) 90.83 (+1.77)
+TACO-RL - with Bleu 28.81 (+16.84) 57.32 (+12.37) 33.99 (+18.16) 46.90 (+14.82) 91.56 (+2.50)
 0.25 (4x compression)
-LLMLingua-2 - CodeSearchNet 2.26 21.99 4.54 14.78 80.05
-LLMLingua-2 - Py150 2.36 (+0.10) 22.37 (+0.38) 4.40 (-0.14) 14.86 (+0.08) 80.11 (+0.06)
-LLMLingua 2.28 (+0.03) 23.64 (+1.65) 5.78 (+1.23) 16.41 (+1.64) 80.24 (+0.18)
+LLMLingua-2 - CodeSearchNet 8.93 41.07 12.30 28.87 88.31
+LLMLingua-2 - Py150 9.35 (+0.42) 41.62 (+0.55) 12.81 (+0.51) 29.21 (+0.34) 88.45 (+0.14)
+LLMLingua 9.06 (+0.13) 41.51 (+0.44) 13.69 (+1.39) 30.15 (+1.28) 88.34 (+0.02)
 Selective Context - - - - -
-TACO-RL - with Rouge1 6.27 (+4.01) 33.19 (+11.20) 13.37 (+8.83) 24.58 (+9.80) 81.75 (+1.70)
-TACO-RL - with Bleu 8.18 (+5.92) 35.49 (+13.50) 17.40 (+12.86) 27.43 (+12.65) 82.24 (+2.19)
+TACO-RL - with Rouge1 17.59 (+8.66) 50.87 (+9.80) 23.03 (+10.73) 38.48 (+9.61) 90.22 (+1.91)
+TACO-RL - with Bleu 24.23 (+15.30) 54.89 (+13.82) 31.26 (+18.96) 44.61 (+15.74) 91.14 (+2.83)
 0.20 (5x compression)
-LLMLingua-2 - CodeSearchNet 1.97 21.01 3.99 14.21 79.96
-LLMLingua-2 - Py150 1.88 (-0.09) 21.07 (+0.06) 3.83 (-0.16) 14.14 (-0.07) 79.98 (+0.02)
-LLMLingua 1.45 (-0.52) 21.39 (+0.37) 4.27 (+0.29) 14.74 (+0.53) 79.94 (-0.02)
+LLMLingua-2 - CodeSearchNet 7.31 38.88 10.56 27.21 87.87
+LLMLingua-2 - Py150 7.76 (+0.46) 39.09 (+0.21) 11.00 (+0.45) 27.46 (+0.26) 88.00 (+0.13)
+LLMLingua 6.57 (-0.73) 38.63 (-0.25) 11.13 (+0.58) 27.81 (+0.60) 87.74 (-0.13)
 Selective Context - - - - -
-TACO-RL - with Rouge1 4.98 (+3.01) 31.84 (+10.83) 12.03 (+8.04) 23.68 (+9.47) 81.63 (+1.67)
-TACO-RL - with Bleu 7.02 (+5.05) 34.96 (+13.95) 16.90 (+12.91) 26.90 (+12.69) 82.16 (+2.20)
+TACO-RL - with Rouge1 14.15 (+6.84) 47.94 (+9.06) 20.20 (+9.65) 36.10 (+8.89) 89.76 (+1.89)
+TACO-RL - with Bleu 21.13 (+13.82) 52.67 (+13.79) 29.14 (+18.58) 42.84 (+15.63) 90.81 (+2.94)
 0.166 (6x compression)
-LLMLingua-2 - CodeSearchNet 1.64 20.14 3.48 13.56 79.87
-LLMLingua-2 - Py150 1.78 (+0.14) 20.69 (+0.55) 3.47 (-0.01) 13.95 (+0.39) 79.92 (+0.05)
-LLMLingua 1.04 (-0.61) 19.93 (-0.21) 3.42 (-0.06) 13.70 (+0.14) 79.72 (-0.15)
+LLMLingua-2 - CodeSearchNet 6.54 37.28 9.58 26.20 87.64
+LLMLingua-2 - Py150 6.79 (+0.25) 37.63 (+0.35) 9.95 (+0.37) 26.49 (+0.29) 87.69 (+0.05)
+LLMLingua 5.31 (-1.23) 36.62 (-0.66) 9.76 (+0.18) 26.38 (+0.18) 87.36 (-0.29)
 Selective Context - - - - -
-TACO-RL - with Rouge1 4.22 (+2.58) 30.96 (+10.82) 11.58 (+8.10) 23.00 (+9.44) 81.58 (+1.71)
-TACO-RL - with Bleu 6.01 (+4.37) 34.31 (+14.17) 16.10 (+12.62) 26.36 (+12.80) 82.12 (+2.25)
-Results with Original Prompts 9.80 33.83 15.79 25.39 81.69
+TACO-RL - with Rouge1 12.38 (+5.84) 46.07 (+8.79) 18.73 (+9.15) 34.93 (+8.73) 89.53 (+1.89)
+TACO-RL - with Bleu 18.46 (+11.92) 50.99 (+13.71) 27.33 (+17.75) 41.34 (+15.14) 90.52 (+2.88)
+Results with Original Prompts 87.89 92.87 88.69 91.25 98.61
 Table 6: Performance metrics for different models across various compression rates on the CodeSearchNet Dataset. The scores for Selective Context are not added as it struggles to compress code data.
@@
 Values in parentheses indicate deltas from the original LLMLingua-2 baseline.
+Metric Mean Std 95% CI P-value
+BLEU 21.2461 0.0991 [21.1230, 21.3692] 1.0000
+ROUGE1 55.2168 0.0728 [55.1265, 55.3071] 1.0000
+ROUGE2 31.7949 0.0387 [31.7469, 31.8430] 1.0000
+ROUGEL 41.9159 0.0309 [41.8775, 41.9542] 1.0000
+BERTScore 90.9349 0.0060 [90.9274, 90.9424] 1.0000
+Table 8: One-sample t-test results for 2x compression over 5 runs on the MeetingBank Dataset for TACO-RL.
+The results show mean scores, standard deviation, 95% confidence intervals, and p-values for various evaluation metrics.
+Method BLEU ROUGE1 ROUGE2 ROUGEL BERTScore
+LLMlingua-2 - Wikitext 16.71 52.58 27.73 39.05 90.47
+Original Prompts 21.50 55.19 33.03 42.90 91.12
+TACO-RL - with entropy term 21.35 55.34 31.88 42.17 90.95
+TACO-RL - without entropy term 17.62 53.39 28.59 39.49 90.57
+Table 9: Performance comparison of TACO-RL with and without the entropy term $\lambda H(\mathbf{p})$ on MeetingBank dataset at 2x compression ratio. Best results are shown in bold.